OpenStudy (perl):

how will f (-1 - x ) look relative to the graph of f(x) . my list of transformations is f( x -1) , f(-x), and such, i dont have f ( -x - 1 )

OpenStudy (amorfide):

rearrange into a format that is better of known by this i mean f(-1-x) same as f(-x-1) so you know by the -1 it means it is translated 1 unit in the positive x direction, and the -x means it is reflected someone correct me if I'm wrong

OpenStudy (anonymous):

-1 - x = -(x+1)

OpenStudy (perl):

i have this website http://www.geom.uiuc.edu/docs/reference/CRC-formulas/node3.html

OpenStudy (perl):

f ( -x - 1) = f ( - ( x + 1) )

OpenStudy (perl):

so if i had f(x) = x^3, what will it look like , f( -1 - x )